An electrical load operates at 240 V rms. The load absorbs an average of 8 KW at lagging power factor of 0.6 The reactive power Q is:


 Select one:

а. 6.48L25.84°Ω

b. 5.04L45.57°Ω

с. 4.32L53.13°Ω

d. 5.76L36.87°Ω

е. З.бL60°Ω


1
Expert's answer
2021-06-09T14:13:02-0400
"\cos ϕ = 0.6\\\phi=53.13\degree\\Q=\frac{8}{0.6}\sin{53.13}=10.67\text{ kVAR}"